Atoms of CO2 decay from higher to lower level with a spread of frequency 10kHZ. what will be the line shape if broadening mechanism is found Doppler mechanism.
1
Expert's answer
2020-05-25T11:19:40-0400

The line shape will be Gaussian:


"G(\omega)=\frac{1}{\sqrt\pi\Gamma}\text{exp}\bigg[-\frac{(\omega-\omega)^2}{\Gamma^2}\bigg]."